OpenStudy (anonymous):

Express the complex number in trigonometric form. -5i

OpenStudy (kainui):

Show your best guess and I'll help you along.

OpenStudy (kainui):

\[re^{i \theta}=r ( \cos \theta + i \sin \theta)\]

OpenStudy (anonymous):

ok 5(cos 180° + i sin 180°)

OpenStudy (kainui):

No, because if you calculate cos(180)=-1 and sin(180)=0

OpenStudy (kainui):

So you'll get -5 instead of -5i which is what you want. You need to have the cosine (real) part be zero and the sine(imaginary) part be -1.